# Towards a turnkey approach to unbiased Monte Carlo estimation of smooth functions of expectations

### v2

@@ B.4.1. Second moment of cycling variables
where $\beta$, defined in (10), is the square-root of the second moment of $\widetilde{X}_{i}$. Assumption 3-(i) assumes $|\gamma_{k}|$ is uniformly bounded (by some constant $c$) while condition (10) assumes $\beta<1$. Hence $\sum|\gamma_{k}|a_{k}$ (of Assumption 1) is bounded above by $\sum c\beta_{k}<\infty$. Thus, Lemma 2 applies to $\hat{f}=\hat{f}^{\mathrm{C}}$.
@@
We only prove case (a). The proof for the second and third case follows the same lines.
@@
The expansion has been organised in three separate rows for the following reason. (Note that each row is non-empty since $r\geq k+l.$) (i) We commence by calculating $\mathbb{E}\left[U_{r,l}^{\mathrm{C}}\prod_{j=1}^{k}\widetilde{X}_{j}\right]$. The first and third row involve non-independent products with the term $\prod_{j=1}^{k}\widetilde{X}_{j}$, whereas for the middle row, the expected value is the product of the expectations. (ii) The previous calculation is sufficient since, by symmetry, $\mathbb{E}\left[U_{r,k}^{\mathrm{C}}U_{r,l}^{\mathrm{C}}\right]=\mathbb{E}\left[U_{r,l}^{\mathrm{C}}\prod_{j=1}^{k}\widetilde{X}_{j}\right]$.
@@ B.4.2. Bound on second moments of cycling variables
The proofs relies on bounding each one of the second moments obtained in Lemma 12 separately. Recall that $\operatorname{cov}\left[U_{r,k}^{\mathrm{C}}U_{r,l}^{\mathrm{C}}\right]=\mathbb{E}\left[U_{r,k}^{\mathrm{C}}U_{r,l}^{\mathrm{C}}\right]-\mathbb{E}\left[U_{r,k}^{\mathrm{C}}\right]\mathbb{E}\left[U_{r,l}^{\mathrm{C}}\right]$. Using the expression for $\mathbb{E}\left[U_{r,k}^{\mathrm{C}}U_{r,l}^{\mathrm{C}}\right]$ in Lemma  12(a), since $\rho>1$, we have
@@
Similarly, for case (b) we find
@@
Finally, case (c) follows straightforwardly
@@
Collecting all the stated, for cases  (a),  (b) and  (c), completes the proof. ∎
@@ B.5. Proof of Proposition 11
    Example 16 shows $\lim_{p\rightarrow 0}\mathbb{E}\left[\operatorname{var}[\widehat{\nabla_{\theta}f}^{\mathrm{S}}|R]\right]>0$ for $f(\cdot)=\log(\cdot)$. Section B.5.1 collects the necessary supporting results to enable this calculation.
@@ E.1.2. High dimensions
[…]
E.2. Independent component analysis
@@
For SGD we use mini-batches of size $2$, while for SAEM we scan the whole data set and, as suggested in Allassonnière and Younes, (2012, Section 3.3), randomly update one component of the latent vector $z$ for each data point $y$, this can be seen as a form of mini-batch sampling as noticed in Kuhn et al., (2020).
@@
E.3. Exponential random graphs
For a given $\theta$, we obtain an unbiased estimate of $\mathcal{Z}(\theta)$ by running a waste-free SMC sampler (Dau and Chopin,, 2022) based on adaptive tempering, i.e. the intermediate distributions are $\pi_{t}(y)\propto\exp\left\{\lambda_{t}s(y)\right\}$, with $0=\lambda_{0}<\ldots<\lambda_{T}=\theta$, and at beginning of iteration $t$, $\lambda_{t}$ is set so that ESS (effective sample size) equals $N/2$, where $N$ is the number of particles. Since we use a waste-free sampler, $N=M\times P$, where $M$ is the number of resampled particles, here $M=100$, and $P$ is the length of the simulated chains, here $P=500$. (For the low variance setting, we use the same approach except $M=10^{3}$.) Note that the initial distribution $\pi_{0}$ is easy to sample from, as it is the uniform distribution over the $\binom{n}{2}$ possible networks. The Markov kernel used at time $t$ to move the particles is a Metropolis kernel based on the following proposal: one edge $(i,j)$ is chosen at random, and $y_{ij}$ is flipped. We use the Python package particles (https://github.com/nchopin/particles) to implement this SMC sampler.

### v3

@@ B.4.1. Second moment of cycling variables
where $\beta$, defined in (10), is the square-root of the second moment of $\widetilde{X}_{i}$. Assumption 3-(i) assumes $|\gamma_{k}|$ is uniformly bounded by some constant $c$ while condition (10) assumes $\beta<1$. Hence $\sum|\gamma_{k}|a_{k}$ (of Assumption 1) is bounded above by $\sum c\beta_{k}<\infty$. Thus, Lemma 2 applies to $\hat{f}=\hat{f}^{\mathrm{C}}$.
@@
We only prove case 1. The proof for the second and third case follows the same lines.
@@
The expansion has been organised in three separate rows for the following reason. Note that each row is non-empty since $r\geq k+l.$ (i) We commence by calculating $\mathbb{E}\left[U_{r,l}^{\mathrm{C}}\prod_{j=1}^{k}\widetilde{X}_{j}\right]$. The first and third row involve non-independent products with the term $\prod_{j=1}^{k}\widetilde{X}_{j}$, whereas for the middle row, the expected value is the product of the expectations. (ii) The previous calculation is sufficient since, by symmetry, $\mathbb{E}\left[U_{r,k}^{\mathrm{C}}U_{r,l}^{\mathrm{C}}\right]=\mathbb{E}\left[U_{r,l}^{\mathrm{C}}\prod_{j=1}^{k}\widetilde{X}_{j}\right]$.
@@ B.4.2. Bound on second moments of cycling variables
The proofs relies on bounding each one of the second moments obtained in Lemma 12 separately. Recall that $\operatorname{cov}\left[U_{r,k}^{\mathrm{C}}U_{r,l}^{\mathrm{C}}\right]=\mathbb{E}\left[U_{r,k}^{\mathrm{C}}U_{r,l}^{\mathrm{C}}\right]-\mathbb{E}\left[U_{r,k}^{\mathrm{C}}\right]\mathbb{E}\left[U_{r,l}^{\mathrm{C}}\right]$. Using the expression for $\mathbb{E}\left[U_{r,k}^{\mathrm{C}}U_{r,l}^{\mathrm{C}}\right]$ in Lemma  121, since $\rho>1$, we have
@@
Similarly, for case 2 we find
@@
Finally, case 3 follows straightforwardly
@@
Collecting all the stated, for cases  1,  2 and  3, completes the proof. ∎
@@ B.5. Proof of Proposition 11
    Example 16 shows $\lim_{p\rightarrow 0}\mathbb{E}\left[\operatorname{var}[\widehat{\nabla_{\theta}f}^{\mathrm{S}}|R]\right]>0$ for $f(\cdot)=\log(\cdot)$. The necessary supporting results to enable this calculation are presented just before this example.
@@ E.2. Toy LVM: High dimensions
E.3. Independent component analysis
@@
For SGD we use mini-batches of size $2$, while for SAEM we scan the whole data set and, as suggested in Allassonnière and Younes, (2012, Section 3.3), randomly update one component of the latent vector $z$ for each data point $y$. This can be seen as a form of mini-batch sampling as noticed in Kuhn et al., (2020).
@@
E.4. Exponential random graphs
For a given $\theta$, we obtain an unbiased estimate of $\mathcal{Z}(\theta)$ by running a waste-free SMC sampler (Dau and Chopin,, 2022) based on adaptive tempering, i.e. the intermediate distributions are $\pi_{t}(y)\propto\exp\left\{\lambda_{t}s(y)\right\}$, with $0=\lambda_{0}<\ldots<\lambda_{T}=\theta$, and at beginning of iteration $t$, $\lambda_{t}$ is set so that ESS (effective sample size) equals $N/2$, where $N$ is the number of particles. Since we use a waste-free sampler, $N=M\times P$, where $M$ is the number of resampled particles, here $M=100$, and $P$ is the length of the simulated chains, here $P=500$. For the low variance setting, we use the same approach except $M=10^{3}$. Note that the initial distribution $\pi_{0}$ is easy to sample from, as it is the uniform distribution over the $\binom{n}{2}$ possible networks. The Markov kernel used at time $t$ to move the particles is a Metropolis kernel based on the following proposal: one edge $(i,j)$ is chosen at random, and $y_{ij}$ is flipped. We use the Python package particles (https://github.com/nchopin/particles) to implement this SMC sampler.
@@ E.4. Exponential random graphs
E.5. Exponential random graphs: Comparison with the exchange algorithm
We compare our approach with one of the state of the art for Bayesian inference on ERG models, i.e. the exchange algorithm of Caimo and Friel, (2011) implemented in the bergm package in R. We use the same setup described in Section 4.5 therein. The shape of the posterior distribution seems to be in good agreement with the obtained in our experiments (see Figure 10).
Refer to captionRefer to captionRefer to captionSimpleCyclingExchange algorithm
Figure 10. Bivariate weighted histograms approximating the posterior distributions obtained with the simple estimator, the cycling estimator and the bergm package implementing the exchange algorithm of Caimo and Friel, (2011).
As we now discuss, there are good reasons in favour of our method. First, the exchange algorithm of Caimo and Friel, (2011) requires running nested MCMC chains where the inner one samples from the likelihood $p(y|\theta)$; in addition, in the version of the algorithm implemented in bergm, population Monte Carlo is employed to improve mixing. As a consequence, the resulting method is rather expensive and, due to the serial nature of MCMC, harder to parallelise than our proposed approach, in which each $\mathcal{Z}(\theta)$ is obtained from different independent runs of SMC which can run in parallel.
Secondly, a fundamental task in the context of network model is the ability to select the model that best fits the data. This type of inference requires estimation of the model evidence $p(y)$. To the best of our knowledge, given the output of the exchange algorithm one can compute the model evidence
$p(y)=p(\theta)\frac{\exp\{\theta^{T}s(y)\}}{\mathcal{Z}(\theta)}\frac{1}{p(\theta|y)}$
in two ways:
  1. (1)
    An approach similar to ours in which $\log\mathcal{Z}(\theta)$ is estimated via path sampling, and $p(\theta|y)$ via the exchange algorithm and a kernel density estimate (KDE) (Caimo and Friel,, 2013, Section 4).
  2. (2)
    Pseudo-likelihood approaches in which the intractable likelihood is replaced by a tractable approximation (Bouranis et al.,, 2018).
The first approach requires running the exchange algorithm to sample from $p(\theta|y)$, from which a KDE is formed. Path sampling with auxiliary MCMC chains is used to sample from $p(y|\theta)$. This approach is clearly only suitable for a small number of parameters due to the instability of the KDE in high dimensions. Also, as mentioned by the authors, this strategy incurs two sources of error when estimating $\log\mathcal{Z}(\theta)$: the Monte Carlo error to approximate the expectation w.r.t. $p(y|\theta)$, and the error due to discretising the path integral. The resulting approximation of the model evidence is naturally biased. Our approach allows computation of the model evidence directly from the samples used to estimate the posterior, without requiring additional samples from $p(y|\theta)$ and without the use of KDE. It also provides unbiased estimators which can be used within pseudo-marginal methods for model selection.
The second approach also cannot provide unbiased estimators of the model evidence and the quality of the approximation largely depends on the pseudo-likelihood used. To obtain a likelihood approximation close to the true one, Bouranis et al., (2018) propose several improvements. We tested the evidence estimation in the bergm package on our toy example and found that the Monte Carlo MLE estimator needed to implement their proposed adjustments to the pseudo-likelihood could not converge since the underlying MCMC was not mixing.
E.6. Comparison with biased estimators
To further motivate the need for unbiased estimators, we compare the MLMC estimator of Shi and Cornish, (2021) and our simple and cycling estimators with their biased counterparts.
E.6.1. Toy LVM
We first consider the toy LVM and two biased estimators:
  • the naive estimator $f(\bar{X})$ where $\bar{X}$ is the mean of the $X_{i}$’s;
  • the IWAE estimator $I_{j}$ given below (18);
and their unbiased counterparts: simple/cycling and MLMC, respectively (see Figure 11 and Figure 12).
Figure 11 shows the behaviour of the relative mean absolute deviation $\mathbb{E}[|f(m)-\hat{f}|]/|f(m)|$ and bias as the expected computational cost increases for the toy LVM model with $d=2$.
[…]
Figure 11. Mean absolute deviation and bias over 100 repetitions against cost for 2 biased estimators (IWAE, naive) and 3 unbiased estimators (MLMC, simple, cycling). The vertical dashed lines are the expected costs used in Figure 2 of the revised/original manuscript, namely $E[R]=6$ and $E[R]=96$.
For large enough cost all estimators except the simple one have very small relative mean absolute deviation. This is because the variance of the simple estimator does not converge to zero as the cost increases, as shown in Example 8. However, for small cost the average (out of 100 replicates) bias of both IWAE and the naive estimator is considerably larger than that of the cycling estimator (see Figure 11).
[…]
Figure 12. Distribution of the estimates of $f(m)$ for small ($E[R]=6$) and high ($E[R]=96$) cost. The dashed horizontal line is the true value.
Figure 12 shows the distribution of the 100 replicates for the smallest cost (left-most vertical dashed line) and the highest one. We can see that cycling provides a provably unbiased estimator without increasing the variability in the original biased naive estimator. When the cost is large it is also clear that cycling outperforms simple since the variability of the latter does not decrease to zero as cost increases.
E.6.2. Independent component analysis
We also compare the plug in estimator $f(\bar{X})$ and the IWAE one for the ICA model using the same setup of Table 3 where we set the number of samples for IWAE and the plug in estimator to $10$.
[…]
SAEM $7.00\cdot 10^{-7}$ 0.22 33
MLMC $2.95\cdot 10^{-6}$ 0.29 11
simple $2.97\cdot 10^{-6}$ 0.27 10
cycling $6.80\cdot 10^{-5}$ 0.24 14
IWAE $3.55\cdot 10^{-6}$ 0.30 7
Plug in $3.62\cdot 10^{-6}$ 0.27 7
E.6.3. Exponential random graphs
[…]
Simple Cycling Plug in
[…]
Moderate 61 0.32 24 0.15 26 0.29
Low 83 0.04 26 0.01 105 0.09
Table 4. Work-normalised variance and mean squared error for $10^{3}$ replicates of $\widehat{\mathcal{Z}(\theta)^{-1}}$ for the simple, the cycling estimator and the plug in estimator.